OpenStudy (anonymous):

A line contains points (2, -2) and (-4, 1). Another line graphed in the same coordinate plane contains points (1, 2) and (-3, 4). Based on the slopes of these lines, are they parallel, perpendicular or neither?

OpenStudy (anonymous):

m = ( 1 +2) / ( -4 -2) = - 3/ 6 = -1/2 m' = ( 4-2) / (-3 -1) = 2/-4 = -1/2 They have same slope - 1/2, they are parallel.